Part A CaO can be used as a drying agent. One such application occurs when water is added to dry concrete or cement. The reaction that occurs is CaO(s)+H2O(l)⇌Ca(OH)2(s) The product is commonly called slaked lime. Assuming the commonly used standard-state temperature of 25∘C, calculate ΔSuniv for this reaction using table from the table below. Substance S∘ [J/(K⋅mol)] ΔH∘f (kJ/mol) CaO(s) 39.9 −635.1 H2O(l) 69.9 −285.8 Ca(OH)2(s) 83.4 −986.1 Express your answer to three significant figures and include the appropriate units. Part B A particular solvent with ΔS∘vap=112.9J/(K⋅mol) and ΔH∘vap=38.0kJ/mol is being considered for an experiment. In the experiment, which is to be run at 75 ∘C, the solvent must not boil. Based on the overall entropy change associated with the vaporization reaction, would this solvent be suitable and why or why not? The solvent would A particular solvent with and is being considered for an experiment. In the experiment, which is to be run at 75 , the solvent must not boil. Based on the overall entropy change associated with the vaporization reaction, would this solvent be suitable and why or why not?1 be suitable because ΔSuniv>0. 2 not be suitable because ΔSuniv>0. 3 be suitable because ΔSuniv<0. 4 not be suitable because ΔSuniv<0.
Answer #1

Express your answer to three significant figures and include the appropriate units

Substance.....S in J/(K*mol)......Delta H f in kJ/mol
CaO (s).......39.9.................... -635.1
H2O(l).........69.9...................... -285.8
Ca(OH)2(s)..83.4.................... -986.1

CaO(s) + H2O(l) <-->Ca(OH)2(s)

Tsurr = 0 deg. C = 273 K

ΔS_surr = -ΔH/Tsurr = : -(-986.1 - 635.1 - 285.8)/298 = 218.8

ΔS_surr = 218.8 J/K

ΔS_total = 218.8 - 26.4
ΔS_total = 192.4 J/K =

particular solvent with Delta S* vap = 112.9 J/(K * mol) and Delta H* vap = 38.0 kJ/mol is being considered for an experiment. In the experiment, which is to be run at 75 C, the solvent must not boil. Based on the overall entropy change associated with the vaporization reaction, would this solvent be suitable and why or why not?

The Solvent Would :

be suitable because ΔS univ>0 .

not be suitable because ΔS univ>0 .-answer

be suitable because ΔS univ<0 .

not be suitable because ΔS univ<0 .

Suniverse=∆Ssystem+∆Ssurroundings

∆ Suniverse=∆Ssystem −∆ H system/T

∆ Suniverse=112.9 JK mol − 38.0kJ/mol/ (75℃+273)×1000J1kJ

∆ Suniverse=112.9 J/K mol−109.19J/K mol

=3.71 JK mol

bp = ∆Hvap / ∆Svap = (38000 J / mole) / (112.0 kJ/ moleK) = 337 K = 64C solvent would boil and would therefore NOT be suitable b/c ΔStotal > 0